OpenStudy (anonymous):

4a=4-b 4a=6-b Use the method of substitution to solve the system of linear equations

Parth (parthkohli):

\[4a= 6-b \\4a =4- b\]That just means this:\[4- b = 6 - b\]

Parth (parthkohli):

What do you think?

OpenStudy (anonymous):

I think the question is leading me to come up with an ordered pair.